3) Compute the Fourier transform of 1 +e-314(20(t) – 1).
Answer #1

Answers Fougies faransfoam of it e -3(1) (Qu(t) - -] Take. qu(t)-1. Sgn (+) fault) -t 7 San(t) -t -1 sgn (t) = u(t) - (-)| + e e-31- alte -3(1) -30 UCE)-e [u(t)-41-6) u(-1) -3(1) Pu(t) foceries transfam Exist only. when tis poistive e 0-314) u(-4Area cindes S(w)= 20(t) शा Ao Ao ॥ 4. 8(6) Ao Ž 24 to Sco) If we put Ao=1 1211.1 f(w). FT (1) 20 8(w) -3.t- Xı(w) - 5 أمال -X₂ (W) 2+3+46-t)ezteedt. jest , (3-jo) dt 0 - e و (سال -3) همان -3 3-uw no 1 Xq (0) 3-jw -3t Ité UCE) est u (-E). 1 į n 8(w)I hope this information will be helpful to you

Please like

Thank you